# Optimal Error Analysis of Channel Estimation for IRS-assisted MIMO Systems

### v1

@@ II Single-Hop IRS-Assisted MIMO System
Suppose each element of the pilot signal matrix $\bm{X}$ is a complex-valued subgaussian random variable  [^2] Note that a complex-valued random variable $X$ is subgaussian if and only if its both real part $\operatorname{Re}\left\{X\right\}$ and imaginary part $\operatorname{Im}\left\{X\right\}$ are real subgaussian random variables. Here are some classical examples of subgaussian distributions. (Gaussian) A standard complex Gaussian random variable $X=\operatorname{Re}\left\{X\right\}+i\operatorname{Im}\left\{X\right\}$ with $\operatorname{Re}\left\{X\right\}$ and $\operatorname{Im}\left\{X\right\}$ being independent and following $\mathcal{N}(0,\frac{1}{2})$, is a subgaussian random variable. (Bernoulli) A Bernoulli random variable $X$ that takes values $-1$ and $1$ with equal probability is a subgaussian random variable. (Digital modulation signals) The phase shift keying (PSK) [60], quadrature amplitude modulation (QAM) [61] and pulse amplitude modulation (PAM) [62] exhibit subgaussian random variable properties. . Let $\delta_{UM}\in(0,1)$ be a positive constant. Then, for any channel tensor $\mathcal{B}\in\mathbb{B}_{N,\mathcal{S}}$, when the number of time slots satisfies
@@
with probability $1-e^{-cT}$, $\mathcal{X}$ satisfies the $UM$-RIP:
where $c$ and $C$ are positive constants.
The proof is presented in Appendix B. Theorem 1 guarantees the RIP for subgaussian measurement ensembles, with the number of time slots $T$ scaling linearly only in relation to the total number of antennas ($UM$) in the UTs  [^3] In a downlink communication scenario, the parameter $T$ will be connected to the total number of antennas ($LP$) in the BSs.. Subgaussian random variables, including Gaussian, Bernoulli, PSK, QAM, and PAM signals, can be conveniently designed as pilot signals in practical applications. Notably, while each element of the Haar-distributed random unitary matrix is a subgaussian random variable, the same does not hold for the elements of the normalized DFT matrix, $\frac{1}{\sqrt{T}}\bm{X}$, a commonly employed pilot signal matrix. Nevertheless, the DFT matrix satisfies (10) when $T\geq UM$, rendering the subsequent conclusions with $\delta_{UM}=0$ applicable to them as well. Our analysis centers on the subgaussian case to streamline the discussion. When RIP holds, then for any two distinct TT format channel tensors $\mathcal{B}_{1},\mathcal{B}_{2}\in\mathbb{B}_{N,\mathcal{S}}$, noting that $\mathcal{B}_{1}-\mathcal{B}_{2}\in\mathbb{B}_{2N,\mathcal{S}}$ is also a TT format channel tensor according to (A) in Appendix A, we have distinct measurements since
[…]
@@
Next, we present a formal analysis of the upper bound for ^$\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}$ with ^$\widehat{\mathcal{B}}\in\mathbb{B}_{N,\mathcal{S}}$. Here, we provide a concise summary of the main result.
@@
(Upper bound of ^$\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}$) Given a channel tensor $\mathcal{B}^{\star}$ in (8) where $\bm{H}$ and $\bm{G}$ are full rank, when each element of the pilot signal matrix $\bm{X}$ is a complex-valued subgaussian random variable and each element in $\mathcal{W}$ follows the complex normal distribution $\mathcal{C}\mathcal{N}(0,\gamma^{2})$, with probability $1-2e^{-c_{1}(LPN+UMN)}$ for a positive constant $c_{1}$, we have
^$\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}\leq O\bigg{(}\frac{\gamma\sqrt{(1+\delta_{UM})(LPN+UMN)}}{(1-\delta_{UM})\sqrt{T}}\bigg{)},$ (12)
@@
The detailed analysis is provided in Appendix C. Theorem 2 guarantees a stable recovery of the ground truth $\mathcal{B}^{\star}$ when the number of time slots $T$ is linearly proportionate to $LPN+UMN$. Combining the condition $T\geq\Omega(UM/\delta_{UM}^{2})$ in (9) with ^$\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}\leq\epsilon$, the number of time slots $T$ should satisfy
$T\geq\max\bigg{\{}\Omega\bigg{(}\frac{UM}{\delta_{UM}^{2}}\bigg{)},\Omega\bigg{(}\frac{(1+\delta_{UM})(LPN+UMN)\gamma^{2}}{(1-\delta_{UM})^{2}\epsilon^{2}}\bigg{)}\bigg{\}}$
to ensure a small recovery error.
@@
Next we consider two special cases: $(i)$ Theorem 2 assumes that the channel matrices $\bm{H}$ and $\bm{G}$ are full rank, but in some scenarios, such as millimeter-wave MIMO systems, the presence of a large number of transmit/receive antennas combined with scattering-poor propagation may lead to low-rank channel matrices $\bm{H}$ and $\bm{G}$ [63, 15]. When the signal travels between the BS and IRS across $r_{1}$ clusters, and between the IRS and the UT across $r_{2}$ clusters, it is reasonable to assume that $\bm{H}$ is a rank-$r_{1}$ matrix and $\bm{G}$ is a rank-$r_{2}$ matrix. By leveraging the covering number of low-rank matrices as introduced in [64, Lemma 3.1] and applying it to (40) in Appendix C, we can readily extend (12) to
^$\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}$
$\leq O\bigg{(}\frac{\gamma\sqrt{(1+\delta_{UM})((LP+N)r_{1}+(UM+N)r_{2})}}{(1-\delta_{UM})\sqrt{T}}\bigg{)},$ (13)
@@
where ^,$\widehat{\mathcal{B}},\mathcal{B}^{\star}\in\{\mathcal{B}=[\bm{H},\mathcal{S},\bm{G}]:\bm{H}\in\mathbb{C}^{LP\times N}\ \text{with}\ \text{rank}(\bm{H})$
$=r_{1},\bm{G}\in\mathbb{C}^{N\times UM}\ \text{with}\ \text{rank}(\bm{G})=r_{2},\text{are unknown and}$
@@
$\mathcal{S}\in\mathbb{C}^{N\times K\times N}\ \text{is known}\}$. It is important to note that a comparison between (12) and (II) reveals that the improvement in recovery error due to low-rankness occurs only when $r_{1}\leq\frac{LPN}{LP+N}$ and $r_{2}\leq\frac{UMN}{UM+N}$.
@@
$(ii)$ It is crucial to emphasize that the elements of the IRS in outdoor scenarios, being exposed to weather and atmospheric conditions, may encounter unknown blockages and time-dependent fluctuations in their phase and amplitude responses [51]. Consequently, the IRS tensor $\mathcal{S}$ deviates from its intended structure, and the assumption of perfect knowledge of all phase shifts at the receiver may not be valid in such cases. To analyze the scenario where $\mathcal{S}$ is unknown, by introducing the covering number for $\mathcal{S}$ to the proof in Appendix C , we can obtain
^$\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}\leq O\big{(}\frac{\gamma\sqrt{(1+\delta_{UM})(LPN+KN+UMN)}}{(1-\delta_{UM})\sqrt{T}}\big{)},$
@@
where ^,$\widehat{\mathcal{B}},\mathcal{B}^{\star}\in\{\mathcal{B}=[\bm{H},\mathcal{S},\bm{G}]:\bm{H}\in\mathbb{C}^{LP\times N},\mathcal{S}\in\mathbb{C}^{N\times K\times N},\bm{G}\in\mathbb{C}^{N\times UM}\ \text{are unknown}\}$. Apart from the additional term $KN$, this upper bound closely aligns with (12). Hence, we omit redundant discussions and concentrate exclusively on the scenario where $\mathcal{S}$ is perfectly known.
Our second goal is to establish a minimax lower bound using Gaussian ensemble design. This lower bound proves to be a valuable tool for gaining insights into the parameters that influence the achievable error of a given problem. Furthermore, it reveals that the upper bound in (12) is tight and optimal. Following the derivation in Appendix D, we arrive at:
@@
(Minimax lower bound of ^$\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}$) In the context of the TT-based ToT regression model in (6), we assume a channel tensor $\mathcal{B}^{\star}$ from (8), where $\bm{H}$ and $\bm{G}$ are full rank and $N\geq C^{\prime}$ with $C^{\prime}$ being a universal constant. Assuming each element of $\bm{X}$ and $\mathcal{W}$ in (6) follows $\mathcal{C}\mathcal{N}(0,1)$ and $\mathcal{C}\mathcal{N}(0,\gamma^{2})$, respectively, we obtain:
$\inf_{\widehat{\mathcal{B}}}\sup_{\mathcal{B}^{\star}\in\mathbb{B}_{N,\mathcal{S}}}\operatorname{\mathbb{E}}{\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}}\geq\Omega\bigg{(}\sqrt{\frac{LPN+UMN}{T}}\gamma\bigg{)}.$ (14)
Note that the constraint $N\geq C^{\prime}$ is not strictly necessary, primarily simplifying the theoretical analysis in Appendix D.

### v2

@@ II-C Optimal error analysis
A linear operator $\mathcal{X}:\mathbb{C}^{LP\times K\times UM}\rightarrow\mathbb{C}^{LP\times K\times T}$ is said to satisfy the $UM$-RIP with a constant $\delta_{UM}$ if
holds for any channel tensor $\mathcal{B}\in\mathbb{B}_{N,\mathcal{S}}$. We now consider two cases:
  • Suppose each element of the pilot signal matrix $\bm{X}$ is an i.i.d. complex-valued subgaussian random variable  [^2] Note that a complex-valued random variable $X$ is subgaussian if and only if its both real part $\operatorname{Re}\left\{X\right\}$ and imaginary part $\operatorname{Im}\left\{X\right\}$ are real subgaussian random variables. Here are some classical examples of subgaussian distributions. (Gaussian) A standard complex Gaussian random variable $X=\operatorname{Re}\left\{X\right\}+i\operatorname{Im}\left\{X\right\}$ with $\operatorname{Re}\left\{X\right\}$ and $\operatorname{Im}\left\{X\right\}$ being independent and following $\mathcal{N}(0,\frac{1}{2})$, is a subgaussian random variable. (Bernoulli) A Bernoulli random variable $X$ that takes values 1-1 and $1$ with equal probability is a subgaussian random variable. (Digital modulation signals) The phase shift keying (PSK) [60], quadrature amplitude modulation (QAM) [61] and pulse amplitude modulation (PAM) [62] exhibit subgaussian random variable properties. Specifically, an i.i.d. sequence of PSK, QAM, or PAM symbols can be generated by first drawing a bitstream as an i.i.d. random process with zero mean and bounded variance, and then mapping each bit independently to the corresponding constellation points. The resulting symbol sequence can thus be treated as an i.i.d. complex-valued subgaussian process with a finite subgaussian norm. . For a positive constant $\delta_{UM}\in(0,1)$, when the number of time slots satisfies
@@
    with probability $1-e^{-cT}$, $\mathcal{X}$ satisfies the $UM$-RIP with positive constants $c$ and $C$.
  • Suppose $\frac{1}{\sqrt{T}}\bm{X}$ is the normalized DFT matrix, when
    $T\geq UM$ (11)
    is satisfied, we have $\frac{1}{T}\|\mathcal{X}(\mathcal{B})\|_{F}^{2}=\|\mathcal{B}\|_{F}^{2}$ with $\delta_{UM}=0$.
The proof is presented in the Appendix C. Theorem 1 guarantees the RIP for subgaussian measurement ensembles and DFT signals, with the number of time slots $T$ scaling linearly only in relation to the total number of antennas ($UM$) in the UTs. When RIP holds, then for any two distinct TT format channel tensors $\mathcal{B}_{1},\mathcal{B}_{2}\in\mathbb{B}_{N,\mathcal{S}}$, noting that $\mathcal{B}_{1}-\mathcal{B}_{2}\in\mathbb{B}_{2N,\mathcal{S}}$ is also a TT format channel tensor according to (2) in the Appendix A, we have distinct measurements since
$\frac{1}{T}\|\mathcal{X}(\mathcal{B}_{1})-\mathcal{X}(\mathcal{B}_{2})\|_{F}^{2}$ $\!\!\!\!=\!\!\!\!$ $\frac{1}{T}\|\mathcal{X}(\mathcal{B}_{1}-\mathcal{B}_{2})\|_{F}^{2}$ (12)
$\!\!\!\!\geq\!\!\!\!$ $(1-\delta_{UM})\|\mathcal{B}_{1}-\mathcal{B}_{2}\|_{F}^{2},$
@@
Next, we present a formal analysis of the upper bound for $\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}$ with $\widehat{\mathcal{B}}\in\mathbb{B}_{N,\mathcal{S}}$, obtained by minimizing (7), independent of the specific optimization algorithm. Here, we provide a concise summary of the main result.
@@
(Upper bound of $\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}$) Given a channel tensor $\mathcal{B}^{\star}$ in (8) where $\bm{H}$ and $\bm{G}$ are full rank, when $\mathcal{X}$ satisfies the $UM$-RIP and each element in $\mathcal{W}$ follows the complex normal distribution $\mathcal{C}\mathcal{N}(0,\gamma^{2})$, with probability $1-2e^{-c_{1}(LPN+UMN)}$ for a positive constant $c_{1}$, the solution in (7) satisfies
$\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}=O\bigg(\frac{\gamma\sqrt{(1+\delta_{UM})(LPN+UMN)}}{(1-\delta_{UM})\sqrt{T}}\bigg),$ (13)
@@
The detailed analysis is provided in the Appendix D. Theorem 2 guarantees a stable recovery of the ground truth $\mathcal{B}^{\star}$ when the number of time slots $T$ is linearly proportionate to $LPN+UMN$. For subgaussian measurements, combining the condition $T=\Omega(UM/\delta_{UM}^{2})$ in (10) with $\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}\leq\epsilon$, the number of time slots $T$ should satisfy
$T=\max\bigg\{\Omega\bigg(\frac{UM}{\delta_{UM}^{2}}\bigg),\Omega\bigg(\frac{(1+\delta_{UM})(LPN+UMN)\gamma^{2}}{(1-\delta_{UM})^{2}\epsilon^{2}}\bigg)\bigg\}$
to ensure a small recovery error. On the other hand, when using the normalized DFT matrix, to guarantee $\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}\leq\epsilon$, the number of time slots $T$ should satisfy
$T\geq\max\bigg\{UM,c\frac{(LPN+UMN)\gamma^{2}}{\epsilon^{2}}\bigg\},$
where $c$ is a positive constant. We note that a commonly considered IRS-assisted MIMO scenario corresponds to $P=1$, i.e., a single base station, under which all theoretical results in this paper remain valid without any modification of the underlying assumptions. In addition, we emphasize that our result does not require any prior knowledge of the channel matrices $\bm{G}$ and $\bm{H}$ or the IRS configuration $\mathcal{S}$; it only relies on their dimensions. This generality naturally includes special scenarios such as channel hardening [63], which arises when $N$ is sufficiently large.
@@
Next we consider three special cases: $(i)$ Theorem 2 assumes that the channel matrices $\bm{H}$ and $\bm{G}$ are full rank, but in some scenarios, such as millimeter-wave MIMO systems, the presence of a large number of transmit/receive antennas combined with scattering-poor propagation may lead to low-rank channel matrices $\bm{H}$ and $\bm{G}$ [64, 15]. When the signal travels between the BS and IRS across $r_{1}$ clusters, and between the IRS and the UT across $r_{2}$ clusters, it is reasonable to assume that $\bm{H}$ is a rank-$r_{1}$ matrix and $\bm{G}$ is a rank-$r_{2}$ matrix. By leveraging the covering number of low-rank matrices as introduced in [65, Lemma 3.1] and applying it to (12) in Appendix D, we can readily extend (13) to
$\hskip-34.14322pt\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}$ (14)
$\hskip-34.14322pt=O\bigg(\frac{\gamma\sqrt{(1+\delta_{UM})((LP+N)r_{1}+(UM+N)r_{2})}}{(1-\delta_{UM})\sqrt{T}}\bigg),$
@@
where $\widehat{\mathcal{B}},\mathcal{B}^{\star}\in\{\mathcal{B}=[\bm{H},\mathcal{S},\bm{G}]:\bm{H}\in\mathbb{C}^{LP\times N}\ \text{with}\ \text{rank}(\bm{H})$
$r_{1},\bm{G}\in\mathbb{C}^{N\times UM}\ \text{with}\ \text{rank}(\bm{G})=r_{2},\text{are unknown and}$
@@
$\mathcal{S}\in\mathbb{C}^{N\times K\times N}\ \text{is known}\}$. It is important to note that a comparison between (13) and (14) reveals that the improvement in recovery error due to low-rankness occurs only when $r_{1}\leq\frac{LPN}{LP+N}$ and $r_{2}\leq\frac{UMN}{UM+N}$.
$(ii)$ In addition to low-rank structures, channel matrices in mmWave MIMO systems also exhibit sparsity in the angular domain due to the limited number of scattering paths [41, 32]. In this case, $\bm{H}$ and $\bm{G}$ can be decomposed as $\bm{H}=\bm{H}_{1}\bm{\Sigma}_{1}\bm{H}_{2}$ and $\bm{G}=\bm{G}_{1}\bm{\Sigma}_{2}\bm{G}_{2}$, where $\bm{H}_{1}\in\mathbb{C}^{LP\times F_{1}}$, $\bm{H}_{2}\in\mathbb{C}^{F_{1}\times N}$, $\bm{G}_{1}\in\mathbb{C}^{N\times F_{2}}$, and $\bm{G}_{2}\in\mathbb{C}^{F_{2}\times UM}$ are known matrices defined in [41], and $\bm{\Sigma}_{1}\in\mathbb{C}^{F_{1}\times F_{1}}$ and $\bm{\Sigma}_{2}\in\mathbb{C}^{F_{2}\times F_{2}}$ are respectively unknown diagonal matrices containing only $s_{1}$ and $s_{2}$ nonzero entries corresponding to the path gains. When $s_{1}\leq\frac{F_{1}}{2}$ and $s_{2}\leq\frac{F_{2}}{2}$, the sparse channel structure admits an upper bound on the recovery error, given by:
$\hskip-34.14322pt\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}$
$\hskip-34.14322pt=O\bigg(\frac{\gamma\sqrt{(1+\delta_{UM})(s_{1}\log(\frac{F_{1}}{s_{1}})+s_{2}\log(\frac{F_{2}}{s_{2}}))}}{(1-\delta_{UM})\sqrt{T}}\bigg),$ (15)
where $\widehat{\mathcal{B}},\mathcal{B}^{\star}\in\{\mathcal{B}=[\bm{H}_{1}\bm{\Sigma}_{1}\bm{H}_{2},\mathcal{S},\bm{G}_{1}\bm{\Sigma}_{2}\bm{G}_{2}]:\bm{\Sigma}_{1}\in\mathbb{C}^{F_{1}\times F_{1}},\bm{\Sigma}_{2}\in\mathbb{C}^{F_{2}\times F_{2}},\|\bm{\Sigma}_{1}\|_{0}=s_{1},\|\bm{\Sigma}_{2}\|_{0}=s_{2}\ \text{are unknown and}\ \mathcal{S}\in\mathbb{C}^{N\times K\times N},\bm{H}_{1}\in\mathbb{C}^{LP\times F_{1}},\bm{H}_{2}\in\mathbb{C}^{F_{1}\times N},\bm{G}_{1}\in\mathbb{C}^{N\times F_{2}}\ \text{and}\ \bm{G}_{2}\in\mathbb{C}^{F_{2}\times UM}\ \text{are known}\}$. The detailed proof has been provided in the Appendix E. As shown in (II-C), when the logarithmic factor is ignored, the recovery error scales only with the sparsity levels $s_{1}$ and $s_{2}$, which explains the practical effectiveness of exploiting sparsity in mmWave MIMO systems.
@@
$(iii)$ It is crucial to emphasize that the elements of the IRS in outdoor scenarios, being exposed to weather and atmospheric conditions, may encounter unknown blockages and time-dependent fluctuations in their phase and amplitude responses [42]. Consequently, the IRS tensor $\mathcal{S}$ deviates from its intended structure, and the assumption of perfect knowledge of all phase shifts at the receiver may not be valid in such cases. To analyze the scenario where $\mathcal{S}$ is unknown, by introducing the covering number for $\mathcal{S}$ to the proof in Appendix D, we can obtain
$\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}=O\big(\frac{\gamma\sqrt{(1+\delta_{UM})(LPN+KN+UMN)}}{(1-\delta_{UM})\sqrt{T}}\big),$
@@
where $\widehat{\mathcal{B}},\mathcal{B}^{\star}\in\{\mathcal{B}=[\bm{H},\mathcal{S},\bm{G}]:\bm{H}\in\mathbb{C}^{LP\times N},\mathcal{S}\in\mathbb{C}^{N\times K\times N},\bm{G}\in\mathbb{C}^{N\times UM}\ \text{are unknown}\}$. Apart from the additional term $KN$, this upper bound closely aligns with (13). Hence, we omit redundant discussions and concentrate exclusively on the scenario where $\mathcal{S}$ is perfectly known.
Our second goal is to establish a minimax lower bound using Gaussian ensemble design. This bound serves as a fundamental tool for understanding how the structural properties of a given model govern its achievable recovery error. Notably, this lower bound depends solely on the structure of the model and is independent of any specific loss function or estimator. Following the derivation in Appendix F, we arrive at:
@@
(Minimax lower bound of $\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}$) In the context of the TT-based ToT regression model in (6), we assume a channel tensor $\mathcal{B}^{\star}$ from (8), where $\bm{H}$ and $\bm{G}$ are full rank and $N\geq C^{\prime}$ with $C^{\prime}$ being a universal constant. Assume that the noise $\mathcal{W}$ in (6) follows $\mathcal{C}\mathcal{N}(0,\gamma^{2})$, and that the pilot matrix $\bm{X}$ is chosen as either (i) an i.i.d. $\mathcal{C}\mathcal{N}(0,1)$ matrix or (ii) a normalized DFT matrix $\frac{1}{\sqrt{T}}\bm{X}$. Under these two choices, we obtain
$\inf_{\widehat{\mathcal{B}}}\sup_{\mathcal{B}^{\star}\in\mathbb{B}_{N,\mathcal{S}}}\operatorname{\mathbb{E}}{\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}}=\Omega\bigg(\sqrt{\frac{LPN+UMN}{T}}\gamma\bigg).$ (16)
Note that the constraint $N\geq C^{\prime}$ is not strictly necessary, primarily simplifying the theoretical analysis. By ignoring the constant term associated with $\delta_{UM}$ in (13), we conclude that the upper bound–independent of any specific optimization algorithm–on the recovery error of the constrained least squares minimization is tight and optimal.
TABLE I: Comparison of Theoretical Guarantees: Ours vs. [31].
Condition for $T$ Upper Bound on Error Lower Bound on Error
[31] $K\cdot\min\{T,LP\}\geq N,T\geq UM$ (necessary) N/A $\Omega\big(\sqrt{\frac{LPNUM}{T}}\gamma\big)$
Ours $T=\Omega\big(\frac{UM}{\delta_{UM}^{2}}\big)$ or $T\geq UM$ (sufficient) $O\big(\sqrt{\frac{(1+\delta_{UM})(LPN+UMN)}{(1-\delta_{UM})^{2}T}}\gamma\big)$ $\Omega\big(\sqrt{\frac{LPN+UMN}{T}}\gamma\big)$
II-D Comparison with Prior Work [31]
Similar to our theoretical result, [31] provides a detailed analysis of the number of time slots required to recover the channel matrices from (2), establishing that $K\cdot\min\{T,LP\}\geq N$ and $T\geq UM$ are necessary conditions. However, their analysis primarily concerns a specific bilinear alternating least squares (BALS) algorithm that minimizes a constrained least squares objective. In particular, the condition $K\cdot\min\{T,LP\}\geq N$ ensures the uniqueness of the solution in the CP decomposition, and $T\geq UM$ guarantees that $\bm{X}$ is left-invertible. In contrast, our results in (10) and (11), offer sufficient conditions for exact recovery under constrained least squares minimization in (7), independent of any specific algorithm.
Moreover, [31] derives a specific lower bound for the CRB, which has been widely used as a benchmark for performance evaluation in IRS-assisted MIMO channel estimation [31, 34, 66, 67, 68, 69]. Specifically, by defining $\bm{S}=\begin{bmatrix}\bm{s}(1)\cdots\bm{s}(K)\end{bmatrix}^{\top}\in\mathbb{C}^{K\times N}$, assuming $\frac{1}{T}\bm{X}\bm{X}^{\rm H}={\bf I}_{UM}$, and letting each entry of $\mathcal{W}$ follow the complex normal distribution $\mathcal{C}\mathcal{N}(0,\gamma^{2})$, the following bound holds:
$\operatorname{\mathbb{E}}\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}$ $\!\!\!\!=\!\!\!\!$ $\operatorname{\mathbb{E}}\|(\widehat{\bm{G}}^{\top}*\widehat{\bm{H}})\bm{S}^{\top}-(\bm{G}^{\top}*\bm{H})\bm{S}^{\top}\|_{F}$ (17)
$\!\!\!\!=\!\!\!\!$ $\Omega\big(\sqrt{\frac{LPNUM}{T}}\gamma\big),$
where the inequality follows from [31, eq. (70)]. We note that (2) is inherently nonconvex, involving a total of $LPN+UMN$ degrees of freedom in the unknown channel matrices. In general, the recovery error is expected to scale with the number of degrees of freedom [70, 71, 72, 51, 50, 40]. However, the bound in (17) scales only with $LPNUM$, as it is derived from a relaxed formulation that transforms the original nonconvex model in (2) into a convex one by treating $\bm{G}^{\top}*\bm{H}\in\mathbb{C}^{LPUM\times N}$ as a single unknown variable.As a result, this bound does not accurately reflect the complexity of the original nonconvex problem and is thus not optimal. In contrast, our work provides the first optimal recovery error bound for channel estimation in IRS-assisted MIMO systems that accounts for the full nonconvex structure of the problem. Recall that the minimax lower bound is given by
$\inf_{\widehat{\mathcal{B}}}\sup_{\mathcal{B}^{\star}\in\mathbb{B}_{N,\mathcal{S}}}\operatorname{\mathbb{E}}{\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}}=\Omega\bigg(\sqrt{\frac{LPN+UMN}{T}}\gamma\bigg),$ (18)
which scales with the total degrees of freedom, $LPN+UMN$, in the unknown channel matrices. A detailed comparison is summarized in Table I.